OpenStudy (is3535):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal to the nearest tenth? Data Set 1: {14, 18, 21, 15, 17} Data Set 2: {15, 17, 22, 20, 16}

OpenStudy (is3535):

@iGreen

OpenStudy (is3535):

@Hero

OpenStudy (anonymous):

Recall that\[\text{mean}=\frac{1}{n}\sum_{k=1}^nx_k\]\[\text{MAD}=\frac{1}{n}\sum_{k=1}^n\left|x_k-\text{mean}\right|\]